# Latent Veracity Inference for Identifying Errors in Stepwise Reasoning

### v2

@@ 1 Introduction
We validate our approach on the logical reasoning benchmark ProntoQA (Saparov & He, 2022), the mathematical reasoning task GSM8K (Cobbe et al., 2021), and commonsense reasoning CommonsenseQA (Talmor et al., 2018) using several open-source LMs (Qwen (Bai et al., 2023; Yang et al., 2024a; b), Llama (Touvron et al., 2023a; b; Grattafiori et al., 2024)). Our method yields consistent improvements in verification accuracy over in-context learning baselines, and scales to longer and more complex reasoning chains. We demonstrate the utility of error-identification in downstream tasks by using veracity assignments as feedback for self-correction and self-improvement (Pan et al., 2024).
@@ 2.3 Amortized Veracity Inference (AVI)
Unlike the test-time search method (VS), $Q(V_{z}\mid x,z)$ is trained to predict veracity without conditioning on the answer $Y$. This provides two primary benefits: (1) it enables rapid run-time inference for verifying a CoT $z$ in a zero-shot manner without test-time search and before seeing the correct answer, and (2) it has the potential to improve reasoning by serving as feedback to identify erroneous reasoning steps in need of correction or re-generation. In some tasks (like ProntoQA and CommensenseQA), reasoning steps predicted to be incorrect by AVI can be negated (by inserting a “not” token), yielding a corrected explanation $z^{\prime}$. The correction is expected to be a more accurate reasoning chain, i.e. $P_{\!\textsc{lm}}(y^{*}\mid x\,z^{\prime})>P_{\!\textsc{lm}}(y^{*}\mid x\,z)$. We empirically validate this hypothesis in § 4.4. For tasks where negation can’t be used for correction, a more general strategy involves simply appending the predicted veracity label (“True” or “False”) to the statement.
@@ 3 Related Work
Our work can be viewed as complementary to self-correction and self-improvement methods, proposing a new meethod for identifying errors that can serve as a useful source of feedback for any self-correction/self-improvement method; alternatively, it can be useful on its own for monitoring purposes. In contrast to standard feedback obtained via few-shot prompting of an LM, our proposed methods (VS and AVI) sample from an intractable posterior distribution in an LVM relating stepwise veracity to final answers, and it does not require examples of correct reasoning chains.
@@ 4 Experiments
In this section we evaluate VS emperically, validating our hypothesis that using $P_{\!\textsc{lm}}(v_{z}\;y^{*}\mid x\,z)$ as a proxy reward improves the correctness of the veracity vector (with respect to the ground-truth $v^{*}_{z}$).
@@
First, we use the ProntoQA benchmark (Saparov & He, 2022), because it gives us the ability to synthesize correct reasoning chains ($z^{*}$) synthetically, facilitating the introduction of controlled errors by corrupting specific steps, resulting in an incorrect proof $\tilde{z}$. Key advantages of ProntoQA include binary veracity labels for each logical deduction steps, and the ability to generate proofs from a fictional ontology to isolate the LM’s logical reasoning capability from it’s ability to retrieve memorized facts/trends acquired during training. Moreover, ProntoQA allows for adjusting the reasoning chain length to test our method’s scalability to more complex scenarios. Finally, we use ProntoQA to assess AVI and its impact on answer prediction ($y^{*}$).
@@ 4.1 Evaluating the Accuracy of Veracity Inference
In the GSM8K mathematical reasoning task, VS consistently outperformed baseline methods across all tested LMs. The remaining gap with ground-truth veracity labels is likely due to the distribution of errors induced by our approach of corrupting CoTs with noise.
In cases where a reasoning step can be viewed as a Boolean assertion (ProntoQA and CommonsenseQA), flagging errors is no different from correcting them. In the setting of GSM8K, however, beyond flagging arithmetic errors, one also needs to re-do the flawed calculations in order to correct the answer or identify additional errors, impacting both the baselines and VS.
The importance of the CoT corruption scheme also highlights the need to evaluate VS under other distributions of erroneous reasoning chains. We conduct such an evaluation on ProntoQA in Appendices C.1 and C.2, and find that VS still outperforms baselines, but by only a modest margin in the case where an LM is tasked to identify errors in the reasoning chains produced by the same LM.
Finally, we evaluate VS in settings where veracity is a categorical variable with more than two classes in Appendix C.5.
@@ 4.4 Evaluation of AVI
Inference-time self-improvement methods are related to self-correction, and aim to improve reasoning through resampling and editing, occasionally utilising a distinct feedback module for guiding correction. Our approach is complementary to these self-improvement methods, as it focuses on identifying statements that need correction, and would therefore fit naturally as a feedback module. In Appendix C.3, we illustrate how AVI can be used as the feedback module in the self-improvment method Self-Refine (Madaan et al., 2023), replacing the few-shot prompted LM (corresponding to our baseline Many2Many) that is typically used for this purpose.
@@ 4.5 Sample efficiency and wall-time
As depicted in Figure 4 (left), VS outperforms all baseline methods in terms of sample efficiency, even when limited to a single sample using greedy tree-based initialization. Performance scales favorably, reaching optimal results at around 100 samples. The zero-shot (1-sample) AVI provides Pareto-optimal performance, comparable to the 100-sample iteration of VS. As shown in Figure 4 (right), VS also uses less inference time per sample compared to other baselines. This efficiency stems from the parallelized reward computations using LM likelihoods evaluated on sampled sequences, eliminating the need for sequential token-level decoding that other baselines require.
@@ 5 Conclusion
There are important limitations to our work. The performance of VS can be sensitive to the distribution of errors, and most of our analysis pertains to artificially corrupted CoTs. Preliminary results suggest that these techniques also work for naturally occurring errors (§C.2), but these claims should be tested in a broader range of dedicated experiments. Ultimately, important applications of error-identification (such as self-correction) will require more than flagging mistakes: inferring veracity $v_{z}$ of a CoT does not render the identity $z$ of the reasoning steps unimportant. In the appendix (§C.3) we prototype a way of integrating AVI with self-improvement methods to resample erroneous statements in $z$, but designing more powerful reasoning systems that combine veracity inference with iterative test-time correction schemes remains an important avenue for future work.
@@ A.4 Amortized Veracity Inference (AVI)
  • Optimizer / LR AdamW (Loshchilov & Hutter, 2019), $1\!\times\!10^{-4}$
@@ C.1 Analysis of Robustness to Different Error Distributions
Reasoning errors across reasoning steps exhibit complex dependencies, and won’t necessarily be distributed uniformly. For instance, mistakes can concentrate near the beginning of a CoT (e.g. when premises are misapplied), or toward the end (e.g. when incorrect conclusions are inferred from preceding assertions). To examine robustness under such conditions, we evaluate three perturbation patterns in ProntoQA: front-side (errors injected in the first half of reasoning steps through negating correct statements), uniform, and back-side (second half of statements are negated). Results obtained with Qwen3-8B are shown in Table 4.
@@
C.2 Correcting LM-generated Reasoning Chains with Negation
@@
The main results of this paper focus on controlled experiments using ProntoQA with known ground-truth reasoning chains $z^{*}$ and corresponding error labels $v_{z^{*}}$, allowing for precise evaluation of the accuracy of the inferred $v_{z}$. However, a practical scenario involves an LM generating its own CoT $z$, with naturally occuring errors. We therefore examine whether our method can be effectively applied to this realistic setting. While we cannot easily evaluate the accuracy of the inferred veracity labels $v_{z}$, we can still assess whether our correction method improves the inference of the final answer $y^{*}$ by correcting $z$ into $\tilde{z}$ using the inferred $v_{z}$ as feedback.
Experimental Setup
Our evaluation involved four steps:
  1. 1.
    Data Collection: Reasoning chains $z$ were generated by prompting Qwen-4B on 5,000 ProntoQA questions $x$.
  2. 2.
    Pseudo-labeling: Generated $(x,z)$ pairs and corresponding answers $y^{*}$ were processed by the VS to derive pseudo-labels $\hat{v}_{z}\in\{\texttt{True},\texttt{False}\}^{|z|}$.
  3. 3.
    Training: Another Qwen-4B model was fine-tuned using AVI to learn a distribution $Q(v_{z}\mid x,z)$ that does not depend on $y^{*}$, using hyperparameters consistent with prior experiments.
  4. 4.
    Evaluation: We evaluated reasoning accuracy (on the final answer $y$) on 1000 separate ProntoQA test instances. Chain-of-thought reasoning using Qwen-4B produced initial answers $y$ from chains $z$. Steps in $z$ labeled as False by $\hat{v}_{z}$ were corrected, producing updated chains $\tilde{z}$ and revised answers. Inference used temperature $T=0.01$.
Table 5 shows that standard self-correction using in-context learning (Many2Many) to provide feedback via veracity assignments does not improve performance: the accuracy even drops from $0.712$ to $0.697$. This highlights that applying correction without reliable veracity identification may in fact be harmful. In contrast, our AVI-based correction raises accuracy to $0.730$, showing that once veracity is accurately identified, correction becomes beneficial rather than detrimental.
These results underscore that it is not the correction procedure itself that drives performance gains, but the quality of veracity identification. Without accurate veracity signals, correction may propagate or even amplify errors.
C.3 Improving LM-generated reasoning chains with Self-Refine
Beyond simple correction via negation in ProntoQA, we further evaluate the applicability of our approach in a more realistic setting where negation-based edits are not easily applicable. In particular, we consider CommonsenseQA, where reasoning chains often involve open-ended commonsense statements. In this case, correction naturally integrates into inference-time self-improvement methods such as Self-Refine (Madaan et al., 2023), which iteratively alternates between a feedback model and a refinement model.
We adapt this framework by replacing the baseline feedback model (which is a few-shot prompted LM, similar to our Many2Many basline) with our AVI machine. Given a CoT $z$ generated by Qwen3-4B, the AVI model first predicts veracity labels $v_{z}$ that identify incorrect reasoning steps. The refinement model (a few-shot-prompted Qwen3-4B model) then resamples downstream reasoning steps, conditioned on these veracity labels. This allows AVI to supply explicit supervision on what to correct, while the refinement model handles the problem of how to correct.
For training, we fine-tuned AVI on 10,000 $(x,y^{*},z)$ examples using pseudo-labels derived from VS, following the same setup as in §2.3. Evaluation was conducted on 1,000 test questions from CommonsenseQA, measuring average accuracy across 5 random seeds.
Table 6: Reasoning accuracy on CommonsenseQA with one iteration of Self-Refine. Accuracy is averaged over 5 seeds ($\pm$ standard deviation).
@@ C.3 Improving LM-generated reasoning chains with Self-Refine
Although the absolute gain from integrating AVI into Self-Refine may appear modest (+0.015+0.015 over the raw baseline), it is nearly twice as large as the improvement obtained with a standard Many2Many feedback model (+0.008+0.008), i.e. $0.015/0.008\approx 1.9\times$. This result highlights that identifying what to correct (via AVI) and deciding how to correct (via Self-Refine) are complementary and orthogonal. Together, they have the potential to result in more correct and useful reasoning chains than either component alone.
C.4 Block Metropolis
@@
C.5 Extending to Categorical Variables
@@ C.6 Correlation Analysis
To validate this hypothesis, we measured the correlation between the joint likelihood and veracity accuracy. For each problem we enumerated all possible veracity vectors ($2^{N}$ assignments for N=7N{=}7 reasoning steps) and computed both their joint likelihood $P_{\!\textsc{lm}}(v_{z}\,y^{*}\mid x,z)$ and their Hamming similarity with the ground-truth vector $v_{z}^{*}$. Table 9 reports the average Pearson correlation across 100 randomly selected ProntoQA samples.
Table 9: Pearson correlation between Hamming similarity and joint likelihood across all possible veracity assignments (7 statements per sample). Results are averaged over 100 ProntoQA samples.
Model Qwen3-4B Qwen3-8B
Pearson Correlation 0.56 0.74

### v3

@@ 1 Introduction
We validate our approach on the logical reasoning benchmark ProntoQA (Saparov and He, 2022), the mathematical reasoning task GSM8K (Cobbe et al., 2021), and commonsense reasoning CommonsenseQA (Talmor et al., 2018) using several open-source LMs (Qwen (Bai et al., 2023; Yang et al., 2024a; b), Llama (Touvron et al., 2023a; b; Grattafiori et al., 2024)). Our method yields consistent improvements in verification accuracy over in-context learning baselines, and scales to longer and more complex reasoning chains. We demonstrate the utility of error-identification in downstream tasks by using veracity assignments as feedback for self-correction and self-improvement (Pan et al., 2024).
@@ 2.3 Amortized Veracity Inference (AVI)
Unlike the test-time search method (VS), $Q(V_{z}\mid x,z)$ is trained to predict veracity without conditioning on the answer $Y$. This provides two primary benefits: (1) it enables rapid run-time inference for verifying a CoT $z$ in a zero-shot manner without test-time search and before seeing the correct answer, and (2) it has the potential to improve reasoning by serving as feedback to identify erroneous reasoning steps in need of correction or re-generation. In some tasks (like ProntoQA), reasoning steps predicted to be incorrect by AVI can be negated (by inserting a “not” token), yielding a corrected explanation $z^{\prime}$. The correction is expected to be a more accurate reasoning chain, i.e. $P_{\!\textsc{lm}}(y^{*}\mid x\,z^{\prime})>P_{\!\textsc{lm}}(y^{*}\mid x\,z)$, and we empirically validate this hypothesis in § 4.4. For tasks where negation can’t be used for correction, a more general strategy involves simply appending the predicted veracity label (“True” or “False”) to the statement, and we showcase this approach in our CommonsenseQA experiments.
@@ 3 Related Work
Our work can be viewed as complementary to self-correction and self-improvement methods, proposing a new method for identifying errors that can serve as a useful source of feedback for any self-correction/self-improvement method; alternatively, it can be useful on its own for monitoring purposes. In contrast to standard feedback obtained via zero- or few-shot prompting of an LM, which has been found to be fragile (in part due to prompt sensitivity) and degrade performance in certain cases (Huang et al., 2023), our proposed method samples directly from the posterior distribution of interest in an LVM relating stepwise veracity to final answers. It does not require examples of correct reasoning chains, and the required instruction-prompt is minimal. Decomposing self-correction into verification and refinement sub-tasks has also been proposed in Zhang et al. (2024c), outside the context of LVMs.
Process reward models (PRMs).
PRMs score the quality of intermediate reasoning steps and are widely used in LM post‑training. They differ from our method in two key ways: (i) typically, PRMs require step-level ground-truth labels (e.g. provided by humans (Lightman et al., 2024)); (ii) supervised/reinforcement-learning fine-tuning methods to train PRMs using only outcome-supervision (final‑answer labels) learn stepwise rewards that represent value/advantage and capture instrumental utility rather than stepwise correctness, and can reward useful‑but‑incorrect steps (Uesato et al., 2022; Zha et al., 2025; Yuan et al., 2025; Cui et al., 2025). In contrast, VS and AVI explicitly target step veracity without the need for process supervision, by approximating latent-veracity inference in an LVM.
@@ 4 Experiments
In this section we evaluate VS emperically, [^1] Source code: https://github.com/alstn12088/veracity_inference validating our hypothesis that using $P_{\!\textsc{lm}}(v_{z}\;y^{*}\mid x\,z)$ as a proxy reward improves the correctness of the veracity vector (with respect to the ground-truth $v^{*}_{z}$).
@@
First, we use the ProntoQA benchmark (Saparov and He, 2022), because it gives us the ability to synthesize correct reasoning chains ($z^{*}$) synthetically, facilitating the introduction of controlled errors by corrupting specific steps, resulting in an incorrect proof $\tilde{z}$. Key advantages of ProntoQA include binary veracity labels for each logical deduction steps, and the ability to generate proofs from a fictional ontology to isolate the LM’s logical reasoning capability from it’s ability to retrieve memorized facts/trends acquired during training. Moreover, ProntoQA allows for adjusting the reasoning chain length to test our method’s scalability to more complex scenarios. Finally, we use ProntoQA to assess AVI and its impact on answer prediction ($y^{*}$).
@@ 4.1 Evaluating the Accuracy of Veracity Inference
In the GSM8K mathematical reasoning task, VS consistently outperformed baseline methods across all tested LMs. The remaining gap with ground-truth veracity labels is likely due to the distribution of errors induced by our approach of corrupting CoTs with noise. In cases where a reasoning step can be viewed as a Boolean assertion (ProntoQA and CommonsenseQA), flagging errors is no different from correcting them. In the setting of GSM8K, however, beyond flagging arithmetic errors, one also needs to re-do the flawed calculations in order to correct the answer or identify additional errors, impacting both the baselines and VS.
The importance of the CoT corruption scheme also highlights the need to evaluate VS under other distributions of erroneous reasoning chains. We conduct such an evaluation on ProntoQA and CommonsenseQA in Appendix C.1, and find that VS outperforms baselines in a wider variety of settings. Finally, we evaluate VS in cases where veracity is a categorical variable with more than two classes in Appendix C.4, and assess the sensitivity of VS accuracy to model size in Appendix C.7.
@@ 4.4 Evaluation of AVI
Inference-time self-improvement methods are related to self-correction, and aim to improve reasoning through resampling and editing, occasionally utilising a distinct feedback module for guiding correction. Our approach is complementary to these self-improvement methods, as it focuses on identifying statements that need correction, and would therefore fit naturally as a feedback module. In Appendix C.2, we illustrate how AVI can be used as the feedback module in the self-improvment method Self-Refine (Madaan et al., 2023), replacing the few-shot prompted LM (corresponding to our baseline Many2Many) that is typically used for this purpose. We find that AVI continues to outperform baselines for this type of application, but note that the magnitude of improvement in reasoning performance is smaller than the gain observed in terms of veracity inference accuracy in the preceding experiments. This suggests that verification alone is not the only bottleneck in self-correction/improvement frameworks (Zhang et al., 2024c), and more work is needed to better understand how to make the best use of a stronger verification signal to guide CoT-(re)generation.
@@ 4.5 Sample efficiency and wall-time
As depicted in Figure 4 (left), VS outperforms all baseline methods in terms of sample efficiency, even when limited to a single sample using greedy tree-based initialization. Performance scales favorably, reaching optimal results at around 100 samples. The zero-shot (1-sample) AVI provides Pareto-optimal performance, comparable to the 100-sample iteration of VS. As shown in Figure 4 (right), VS also uses less inference time per sample compared to other baselines. This efficiency stems from the parallelized reward computations using LM likelihoods evaluated on sampled sequences, eliminating the need for sequential token-level decoding that other baselines require. We include a complementary analysis of the inference-time cost of our approach in Appendix C.6.
@@ 5 Conclusion
There are important limitations to our work. The performance of VS can be sensitive to the distribution of errors, and most of our analysis pertains to artificially corrupted CoTs. Preliminary results suggest that these techniques also work for naturally occurring errors (§C.1), but these claims should be tested in a broader range of dedicated experiments. Ultimately, important applications of error-identification (such as self-correction) often require more than flagging mistakes: inferring veracity $v_{z}$ of a CoT does not render the identity $z$ of the reasoning steps unimportant. In the appendix (§C.2) we prototype a way of integrating AVI with self-improvement methods to resample erroneous statements in $z$, but designing more powerful reasoning systems that combine veracity inference with iterative test-time correction schemes remains an important avenue for future work. Our method could also provide a new training signal for reasoning models, for example by using EM to jointly update the generative model with veracity samples obtained with AVI, or by providing a veracity-specific process-level training signal for PRMs in label-scarcity scenarios.
@@ A.4 Amortized Veracity Inference (AVI)
  • Optimizer / LR AdamW (Loshchilov and Hutter, 2019), $1\!\times\!10^{-4}$
@@ C.1 Analysis of Robustness to Different Error Distributions
Reasoning errors exhibit complex dependencies across reasoning steps, and won’t necessarily be distributed uniformly. We therefore evaluate our method under different error distributions using both synthetic CoT corruption methods (§C.1.1) and naturally-occuring errors (§C.1.2).
C.1.1 Different Patterns of Synthetic CoT Corruption
Mistakes can concentrate near the beginning of a CoT (e.g. when premises are misapplied), or toward the end (e.g. when incorrect conclusions are inferred from preceding assertions). To examine robustness under such conditions, we evaluate three perturbation patterns in ProntoQA: front-side (errors injected in the first half of reasoning steps through negating correct statements), uniform, and back-side (second half of statements are negated). Results obtained with Qwen3-8B are shown in Table 4.
@@ C.1.1 Different Patterns of Synthetic CoT Corruption
C.1.2 Naturally Occurring (LM-Generated) Errors
The main results presented in this paper are derived from controlled experiments using synthetically corrupted reasoning chains, facilitating the evaluation of the accuracy of the inferred veracity assignments $v_{z}$. However, a practical scenario involves an LM generating a CoT $z$ on its own, with errors arising naturally as a result of the limitations of the model’s reasoning capabilities. Here, we examine whether our method generalizes to this more realistic setting.
Reasoning chains $z$ were generated using structured-decoding with a Qwen3-4B model for questions $x$ sampled from ProntoQA and CommonsenseQA datasets ($5,000$ samples each). The resulting $(x,z)$ pairs and corresponding answer labels $y^{*}$ were then subjected to VS to obtain predicted veracities $\hat{v}_{z}\in\{\texttt{True},\texttt{False}\}^{|z|}$.
Table 5: Veracity accuracy in naturally generated CoTs (1,000 test samples for each task).
Hamming Similarity Exact Match Accuracy
Method ProntoQA CommonsenseQA ProntoQA CommonsenseQA
Many2Many 0.74 0.59 0.00 0.00
VS 0.92 0.86 0.50 0.59
We evaluated VS on 1,000 test samples (for each dataset) in terms of veracity inference accuracy (Hamming similarity and exact match accuracy), by comparing predicted $\hat{v}_{z}$ to pseudo-ground-truth labels $v_{z}^{*}$ obtained using a GPT-4.1 oracle. The results shown in Table 5 suggest that VS maintains a strong advantage over the in-context learning baseline (Many2Many) under this natural error distribution.
C.2 Self-Correction/Improvement of LM-Generated Reasoning Chains
Following from the experiment detailed in §C.1.2, where VS was applied to naturally-occurring errors, we then sought to evaluate whether AVI can be used as a feedback signal as part of a self-correction/improvement framework for improving reasoning performance. Concretely, we fine-tuned a Qwen3-4B model using AVI on predicted veracities sampled using VS for $10,000$ $(x,z,y^{*})$ triples, to learn a distribution $Q(v_{z}\mid x,z)$ that does not depend on $y^{*}$. The training setup is otherwise identical to the one described in §2.3. Then, we used AVI to predict erroneous steps in $z$ by inferring stepwise veracities $\hat{v}_{z}\sim Q(v_{z}\mid x,z)$.
@@ C.2 Self-Correction/Improvement of LM-Generated Reasoning Chains
In the case of ProntoQA, we corrected steps predicted to be False by negating them, and queried a Qwen3-4B model to predict the answer $y$ conditioned on the corrected CoT.
Table 6 shows that standard self-correction using in-context learning (Many2Many) to provide feedback via veracity assignments does not improve performance in ProntoQA: the accuracy even drops from $0.712$ to $0.697$. This highlights that applying correction without reliable veracity identification may in fact be harmful. In contrast, our AVI-based correction increases accuracy to $0.730$, showing that once veracity is accurately identified, correction becomes beneficial rather than detrimental. Without accurate veracity signals, correction may propagate or even amplify errors.
Next, we moved beyond simple correction via negation by testing the applicability of our approach in a more realistic setting where negation-based edits aren’t a natural way to correct. In particular, we consider CommonsenseQA, where reasoning chains involve open-ended statements in natural language, and correction is more easily conducted as part of inference-time self-improvement methods such as Self-Refine (Madaan et al., 2023), which iteratively alternates between a feedback model and a refinement model. We adapt this framework by replacing the baseline feedback model (which is a few-shot prompted LM, similar to our Many2Many baseline) with our AVI machine. Given a CoT $z$ generated by Qwen3-4B, the AVI model first predicts veracity labels $v_{z}$ that identify incorrect reasoning steps. The refinement model (a few-shot-prompted Qwen3-4B model) then resamples downstream reasoning steps, conditioned on these veracity labels. This allows AVI to supply explicit supervision on what to correct, while the refinement model handles the problem of how to correct.
Table 7: Reasoning accuracy on CommonsenseQA with one iteration of Self-Refine. Accuracy is averaged over 5 random seeds ($\pm$ standard deviation).
@@
The results are shown in Table 7 for 1,000 test questions from CommonsenseQA. The gain from integrating AVI into Self-Refine is nearly twice as large as the improvement obtained with a standard Many2Many feedback model (+0.008+0.008), i.e. $0.015/0.008\approx 1.9\times$, though in absolute terms may appear modest (+0.015+0.015 over the raw baseline). In general, the trend observed in all our experiments is that our method improves the accuracy of error-identification (whether using VS or AVI) by a larger margin over baselines than it improves reasoning accuracy post-correction, hinting at a bottleneck in reasoning capabilities in the underlying model. This reflects the fact that verification may be easier than generation, and that more work is needed to find ways to make better use of the more robust veracity signal provided by VS and AVI in downstream reasoning tasks.
C.3 Block Metropolis
@@
C.4 Extending to Categorical Variables
@@ C.5 Correlation Analysis
To validate this hypothesis, we measured the correlation between the joint likelihood and veracity accuracy. For each problem we enumerated all possible veracity vectors ($2^{N}$ assignments for N=7N{=}7 reasoning steps) and computed both their joint likelihood $P_{\!\textsc{lm}}(v_{z}\,y^{*}\mid x,z)$ and their Hamming similarity with the ground-truth vector $v_{z}^{*}$. Table 10 reports the average Pearson correlation across 100 randomly selected problems drawn from ProntoQA and CommonsenseQA datasets.
Table 10: Pearson correlation between Hamming similarity and joint likelihood across all possible veracity assignments (7 statements per sample). Results are averaged over 100 samples for each dataset.
Model ProntoQA CommonsenseQA
Qwen3-4B 0.56 0.72
Qwen3-8B 0.74 0.78
LLaMA 3B 0.72 0.67
LLaMA 8B 0.70 0.79
@@
C.6 The Cost of Veracity Search on Inference-Time
Table 11: Average VS wall-time over 5 independent runs as a function of CoT length.
Number of Reasoning Steps Time (sec)
3 7.33
5 7.83
7 11.14
9 12.19
11 13.09
In VS, the CoT $z$ is fixed, and we score a proposed veracity assignment $v_{z}$ via the joint likelihood $P_{\mathrm{LM}}(v_{z},y^{*}\mid x,z)$, where $y^{*}$ is the correct answer to the question in $x$. This likelihood is computed with teacher‑forcing: the concatenated sequence $(v_{z},y^{*})$ is processed in a single forward pass (prefill stage), and logits for all tokens are produced under causal self‑attention. This avoids the substantially more expensive sequential autoregressive decoding required by in‑context learning baselines (Many2Many). However, because VS typically requires scoring multiple proposals as part of the search procedure, this advantage diminishes as the number of VS iterations increases. Empirically, we find that our method provides a favorable trade‑off between inference time and veracity accuracy, as shown in Figure 4, but we provide a complementary analysis of the computational cost incurred from VS iterations below.
Let $L_{\text{context}}=|(x,z)|$ and $L_{\text{tail}}=|(v_{z},y^{*})|$. Naively, across $N$ VS iterations, the cost scales as $O\big(N(L_{\text{context}}+L_{\text{tail}})^{2}\big)$ due to the quadratic cost of self-attention. We can reduce computational requirements with prefix key‑value caching for the fixed context $(x,z)$, so the overall cost becomes $O\big(L_{\text{context}}^{2}\big)+O\big(N(L_{\text{context}}\,L_{\text{tail}}+L_{\text{tail}}^{2})\big)$. For proposals that differ from a reference $v_{z}$ at a single position $i$, we can also cache the extended prefix $(x,z,v_{z}[:i])$. If $i$ is uniformly distributed, the expected suffix length is $L_{\text{tail}}/2$, yielding a constant‑factor reduction: $O\big(L_{\text{context}}^{2}\big)+O\big(N(L_{\text{context}}\,L_{\text{tail}}/2+L_{\text{tail}}^{2}/4)\big)$. Finally, batching proposals that share the same cached prefix (or the same divergence index $i$) lets the GPU process many tails in parallel, improving utilization and reducing wall‑clock time. In other words, we can trade memory for faster runtime, if desired.
We empirically evaluate how wall‑clock time varies as a function of CoT length using Qwen3‑4B-based VS in ProntoQA, for CoTs with lengths ranging from 3 to 11 steps. A single GPU (RTX A6000 ADA) was used, and VS involved 100 Metropolis iterations. The measurements displayed in Table 11 reflect end‑to‑end VS, including greedy initialization and simulated annealing.
C.7 Veracity Search on Smaller and Larger Models
The original Veracity Search (VS) experiments were performed on 3B–8B models. To examine model-size sensitivity, we additionally evaluated a smaller model (Qwen3-1.7B) and a larger model (Qwen3-14B). The experimental design and search hyperparameters match those used to produce the results in Table 1.
Table 12: Veracity inference accuracy (Hamming similarity) across three benchmarks (average over 1000 samples each).
GSM8K ProntoQA CommonsenseQA
1.7B 14B 1.7B 14B 1.7B 14B
Many2Many 0.512 0.675 0.529 0.714 0.488 0.524
VS 0.657 0.833 0.713 0.928 0.832 0.980
Across all three tasks, larger models provide more accurate veracity predictions when using either VS or the in-context learning baseline (Many2Many). VS maintains a consistent advantage over this baseline, indicating that the joint-likelihood used to guide VS becomes increasingly informative as model capacity increases.
Large Language Model Usage
Large language models (LLM) were used only for minor polishing of the writing quality both in the main text and in the code. They were also used to assist with debugging. Any modification to the text that was suggested by an LLM to improve clarity was verified by the authors. LLMs were not used to generate the ideas, methods, experimental designs, and analyses of results that are presented in this paper.